# Labels or Input? Rethinking Augmentation in Multimodal Hate Detection

### v1

@@ Introduction
Figure 2 shows the distribution of these quality scores. In 7 out of the 200 cases, the multimodal generation agent fails to overlay the caption onto the image (no caption alignment score is assigned).
We have added a separate error analysis section with demonstrated examples in SI.
@@
The 89% agreement rate confirms the reliability of scaled labels, supporting their use for capturing nuanced hatefulness beyond binary classification. Human evaluations of the counterfactually neutral memes validate our augmentation pipeline, showing high quality across multiple dimensions. While some errors remain, the approach offers a scalable method for generating diverse, bias-reducing training data. This supports broader efforts to build more robust and fair multimodal systems.
@@ Conclusion
In this paper, we present a dual-pronged framework to advance multimodal hate detection through prompt optimization and multimodal data augmentation. We demonstrate that how we frame the task (via prompts) and what data we train on (via counterfactual augmentation) are critical decisions driving robust model performance.
Our prompt optimization experiments—centered on InternVL2—show that prompt structure and label granularity significantly influence outcomes. InternVL2 achieves the highest F1-scores in different settings under both fine-tuning and in zero-shot prompting. These also offer a more context-sensitive foundation for content moderation systems.
Complementing this, we introduce a novel multi-agent pipeline that generates 2,479 counterfactually neutral memes to debias models’ learning to identify hate. This is, to our knowledge, the first large-scale demonstration of counterfactual meme augmentation for hate detection. The extended dataset leads to consistent F1 improvements, particularly in models capable of deeper semantic reasoning. These gains are most pronounced in handling edge cases and class imbalance, affirming the value of data-centric strategies.
Together, our contributions break new ground by offering an integrated framework that combines prompt design with systematic data augmentation for hateful content detection. Future extensions could explore addressing the limitations of this work, such as its focus on text-centric hate, and finding computationally inexpensive alternatives for multimodal data augmentation. The generalizability of our findings is also confined to a single dataset. While we offer an ethical framework for debiasing hate datasets, the ethical implications of using generative models for content modification may require further regulatory scrutiny.
@@ Acknowledgment
This work is supported by the Singapore Ministry of Education AcRF TIER 3 Grant (MOE-MOET32022-0001). We greatly thank Insyirah Mujtahid, Shaz Furniturewala, and Jingwei Gao for their helpful input and assistance during various stages of this work.
Appendix A Supplementary Information
Appendix B Prompt Optimization
Prompting Components
We organize all prompting strategies in this paper into modular components (see Table 4). The simple module poses direct questions to the model. The category module breaks the task into hatefulness subcategories. The scale module requires numerical outputs (e.g., 0–9), while the binary module expects true/false responses. To reduce conversion errors, we apply output constraints within the scale or binary groups. Although some other strategies perform better, we restrict prompts to clean combinations of modules to control for confounders.
Detailed Experimental Settings
The experiments are organized along three dimensions. The first dimension concerns the model category: multi-modal prompting on a large pre-trained model, InternVL2, multi-modal fine-tuning of the same model, and unimodal fine-tuning on a smaller textual model, DistilBERT. To simulate typical computational constraints, we are using LoRA, a low-rank adaptation technique, to reduce the computational overhead during fine-tuning of 8B models.
The second dimension relates to prompting strategies: either a basic prompt designed for the classification task (prompt component: simple) or a more elaborate prompt outlining specific categories of hatefulness (prompt component: category; see Table 4). The third dimension addresses output format: either binary or scaled outputs. The binary label asks for a direct classification (i.e., True or False; label component Binary), while the scale-based format captures more nuanced levels of hatefulness (label component Scale). All component settings are kept consistent across combinations for fair comparison.
As the original dataset does not include scale-based labels, we use GPT-4o-mini as a teacher model to generate them on a sample. To ensure label accuracy, we manually filter entries by cross-checking with the binary ground-truth. For example, instances labeled hateful in the ground-truth but rated as 1 (not hateful) on the 0–9 scale have been discarded. This ensures correctness while enriching the dataset with scale-level and multi-modal information from the teacher model. After this, we end up with approximately 2K instances for training purposes.
For InternVL2, we use its default fine-tuning hyperparameters. For DistilBERT, the following training configuration is applied uniformly: the learning rate is set to $2\times 10^{-5}$, with a batch size of 12 for both training and evaluation per device, trained for 12 epochs. We use a weight decay of 0.01, with evaluation and saving strategies set to occur at every epoch, and enabled loading of the best model at the end.
Appendix C Multimodal Augmentation
Prompting Components
We employ multiple agents in a structured manner for the multimodal data augmentation pipeline. The prompts and generation parameters are summarized in Table 5, with other parameters set to the models’ default values. For generating image descriptions and analyzing hatefulness in background and caption text, we use the open-source models InternLM-XComposer2.5-7B and Qwen2.5-14B, respectively. For generating new non-hateful captions and meme regeneration, we utilize models from OpenAI and Gemini, which are closed-source. These models are selected due to the complexity of the tasks and their state-of-the-art performance.
Detailed Experimental Settings
The experiments are conducted in three main directions: two unimodal (text and vision separately) and one multimodal. We evaluate the effectiveness of the extended dataset with non-hateful augmented memes. For all experiments, state-of-the-art encoder models (base variant) are fine-tuned on two types of training data. As the addition of non-hateful memes increases data imbalance, we apply downsampling, using a random seed of 42 for all processes, including training and checkpoint loading, to ensure reproducibility. Only the last two layers of each encoder are fine-tuned to prevent overfitting. Training is performed for up to 100 epochs with early stopping (patience set to 3, using weighted-F1 score on the validation set as the criterion), a batch size of 128, a learning rate of $2\times 10^{-5}$, and the AdamW optimizer to minimize cross-entropy loss.
Appendix D Error Examples in Augmented Data
To assess the effectiveness of our two data augmentation strategies: (1) scaling binary labels to continuous hatefulness scores and (2) generating new multimodal hate examples, we conducted a comprehensive human evaluation of the resulting datasets.
For the scaled label augmentation, we observe that approximately 11% of the examples yielded human evaluations that disagreed with the predicted hatefulness scores. As shown in Table 6, several of these disagreements stem from indirect or context-dependent hate speech, which can be subtle and difficult for even advanced models like GPT-4o-mini to accurately evaluate. For example, memes involving racial undertones or cultural stereotyping (e.g., Examples 1 and 3) may appear humorous on the surface but carry implicit offensive content. This highlights the challenge of detecting implicit and contextual hate, and suggests that better modeling of sociocultural nuance is needed.
In the case of multimodal data augmentation (Table 7), we identified several quality issues across a small subset of the generated samples:
  • In 7 out of 200 cases, the output image lacks any caption, although the visual background was still appropriate and non-hateful (like Example 1).
  • Example 2 retains a background similar to the original hateful meme, but the mismatch between the visual context and the newly added non-hateful caption results in a low-quality augmentation.
  • Some samples exhibit formatting errors, such as misspellings or unnatural alignment of image and text. This can also be seen in Example 2.
  • In case of Example 3, the background context is lost. The original meme features a person with a disability, while the generated version failed to preserve this detail, resulting in semantic drift.
  • Another sample shows (Example 4) a directional inconsistency in a character’s face, which disrupts the visual coherence and makes the meme appear unnatural.
Dataset Prompt Example
{{Simple}} + {{Binary}} {{Your task is to analyze this given image and its caption to identify if there’s any forms of hateful content.}} {{If it’s hateful, return the ‘TRUE’ — ‘FALSE’.}}
{{Simple}} + {{Scale}} {{Simple}} {{You will be provided with the image, based on this label and the content of the image, you will assign a hatefulness score on a scale from 0 to 9, where: 0  The image is not hateful at all.9  The image is extremely hateful. Return the score as an integer in range 0 to 9.}}
{{Category}} + {{Binary}} {{Simple}}{{Try to focus on the presence of any element that relates to any of the following: 1. Sexual aggression: a. Homophobia and Transphobia: This category encompasses hate speech targeting LGBTQ+ individuals, including slurs, derogatory comments, and rhetoric that seeks to undermine or dehumanize people based on their sexual orientation or gender identity. b. Misogyny and Sexism: This category includes hate speech directed at women or based on gender. It covers derogatory language, stereotypes, and rhetoric that perpetuate gender inequality, objectification, and violence against women. 2. Hate based on ideology: a. Political Hate Speech: This category includes hate speech that is politically motivated, often targeting individuals or groups based on their political beliefs or affiliations. It may include inflammatory language, threats, and rhetoric designed to polarize or incite violence within political contexts. 3. Racism and xenophobia: a. COVID-19 and Xenophobia: This category includes hate speech that arose during the COVID-19 pandemic, often targeting specific ethnic groups or nationalities. It includes xenophobic language blaming certain groups for the spread of the virus, as well as fear-mongering and scapegoating related to the pandemic. b. Racism Against Black People: This category focuses on hate speech directed at Black individuals or communities. It includes racial slurs, stereotypes, dehumanization, and other forms of derogatory language that perpetuate racial discrimination and inequality. c. Racist Hate Against Other Ethnic Groups: This category includes hate speech directed at various ethnic groups other than Black individuals. It covers a range of racial slurs, xenophobic language, dehumanization, and derogatory remarks targeting specific ethnicities or nationalities. d. White Supremacy: This category includes hate speech promoting white supremacist ideology, often intertwined with Christian extremist views. It includes rhetoric that advocates for racial superiority, anti-immigrant sentiments, and the dehumanization of non-white groups, sometimes using religious justifications for these beliefs. 4. Bigotry: a. Anti-Muslim and Islamophobic Hate: This category comprises hate speech aimed at Muslims and Islam. It includes language that promotes fear, hatred, dehumanization, or prejudice against Muslim individuals or communities, often using religious and cultural references to incite hostility. b. Anti-Semitic Hate: This category focuses on hate speech directed at Jewish people and Judaism. It includes references to historical anti-Semitic tropes, conspiracy theories, and other forms of rhetoric that seek to dehumanize or discriminate against Jewish individuals and communities. 5. Miscellaneous Hate Speech: This category captures various forms of hate speech that do not fit neatly into the other categories. It includes a wide range of derogatory language and expressions that target individuals or groups based on different aspects of identity or personal characteristics. This category includes hate speech that targets individuals based on their physical or mental disabilities. It often includes derogatory language that mocks or devalues people with disabilities, promoting harmful stereotypes and exclusion.}} {{Binary}}
{{Category}} + {{Scale}} {{Simple}}{{Category}}{{Scale}}
  • Note. a) Brackets are only used for clarification; b) each color represents a unique component of prompts.
Table 4: Prompts used for prompt optimization framework experiment.
Pipeline Component Model Generation Params* Prompt Used
Background Description
Generation
InternLM-XComposer2.5-7B
do_sample=False,
num_beams=3,
top_p=None,
use_meta=True
Analyze and describe the given image’s visuals in short. Do not include the description about the text written on the image.
Hate Component
Analyser
Qwen2.5-14B max_new_tokens=1 You are a social media content moderator. You are given a text caption of a meme and your task is to identify whether it’s hateful? Answer in ‘Yes’ or ‘No’ only. >>>Image Caption: <TEXT CAPTION> >>>isHateful:
You are a social media content moderator. You are given a text description of the background image of a meme and your task is to identify whether it’s hateful? Answer in ‘Yes’ or ‘No’ only. >>>Image Description: <IMAGE DESCRIPTION> >>>isHateful:
New Non-hateful
Caption Generation
GPT-4o-mini temperature=0 You are a helpful social media expert with deep knowledge of memes and inclusive language. Your task is to rewrite a hateful meme caption so that it becomes non-hateful and bias-reducing, while staying relevant to the background image description. The revised caption must: - Preserve reference to the originally targeted group or person (e.g., Muslims, Women, Immigrant, etc.), if they are mentioned. The new caption should try to retain the word related to target group in a respectful and bias-correcting way. - Transform the message to be neutral or positively framed toward the mentioned group/person. - Help counteract harmful stereotypes or biases. - Remain coherent and contextually appropriate for the image. Strict Rule: Only return the revised caption. Do not include explanations, disclaimers, or any other text. Meme background image description: ‘<IMAGE DESCRIPTION>’ Meme caption text: ‘<TEXT CAPTION>’ Revised non-hateful meme caption:
Meme Image
Regeneration
Gemini-Flash-2.0-exp
temperature=0,
max_output_tokens=100,
response_modalities=[“image”,“text”],
safety_settings=“BLOCK_NONE”,
response_mime_type=“text/plain”
You are a creative social media expert with deep knowledge of internet memes. Your task is to generate a humorous and engaging meme image that is non-hateful. Use the given background description and overlay the provided text in a visually appealing way. Background image description: ‘<IMAGE DESCRIPTION>’ Meme text: ‘<NEW NON-HATEFUL TEXT CAPTION>’ Ensure the meme aligns with popular internet humor while remaining appropriate and widely relatable.
Table 5: Prompts used for multimodal data augmentation pipeline. Note: Any additional parameters are set to their default values.
Example 1 Example 2 Example 3 Example 4
Meme
HS
6 2 6 0
Example 1 Example 2 Example 3 Example 4
Hateful Meme
Non-hateful Meme

### v2

@@ Introduction
Figure 2 shows the distribution of these quality scores. We identified several quality issues across a small subset of the generated samples (examples in Table 7). In 7 out of 200 cases, the output image contains no caption, even though the visual background remains appropriate and non-hateful, as in Example 1. In Example 2, the background closely resembles the original hateful meme, but a mismatch between the visual context and the newly added non-hateful caption leads to a low-quality augmentation; this example also exhibits formatting issues such as misspellings or unnatural alignment between image and text. In Example 3, the background context is not preserved: the original meme depicts a person with a disability, but this detail is lost in the generated version, resulting in semantic drift. Finally, Example 4 shows a directional inconsistency in a character’s facial orientation, which breaks visual coherence and makes the meme appear unnatural.
@@
The 92.5% agreement rate confirms the reliability of scaled labels, supporting their use for capturing nuanced hatefulness beyond binary classification. Human evaluations of the counterfactually neutral memes validate our augmentation pipeline, showing high quality across multiple dimensions. While some errors remain, the approach offers a scalable method for generating diverse, bias-reducing training data. This supports broader efforts to build more robust and fair multimodal systems.
@@ Conclusion
This work advances multimodal hatefulness classification by addressing two core challenges highlighted in computational social science: the need for more disciplined supervision and the need for higher-quality data. Both concerns stem from the broader insight that robust moderation systems depend not only on powerful models, but also on how tasks are framed and how data is constructed. In line with this motivation, we evaluate two complementary strategies—prompt optimization and multimodal data augmentation—that directly respond to the limitations of existing annotation pipelines and the artifacts present in widely used multimodal hate datasets.
Our findings show that prompt design materially shapes downstream performance. Simple prompts suffice for binary detection, while structured category prompts become crucial when labels encode finer degrees of hatefulness. For scaled labels, prompting can even rival full fine-tuning, particularly for large VLMs such as InternVL2. These results highlight that careful prompt construction can serve as a lightweight but powerful mechanism for improving supervision quality in resource-constrained settings.
We also develop a multi-agent augmentation pipeline that produces 2,479 counterfactually neutral memes without introducing new hateful content. This provides an ethically grounded method for refining training distributions and reducing overreliance on spurious text–image correlations. The extended dataset yields consistent improvements in weighted F1 across unimodal and multimodal models and is especially effective for architectures with sufficient capacity to integrate subtle semantic cues. Human evaluations further validate the augmented samples, confirming their reliability as bias-reducing training instances and illustrating the promise of counterfactual data for strengthening semantic robustness.
Taken together, these contributions demonstrate that improving multimodal hate detection requires attention to the quality of task framing and the integrity of training data. We offer a pathway toward moderation systems that are more resilient to dataset artifacts, better aligned with sociocultural nuance, and more responsive to the ethical obligations of studying online harm. Future work should explore scaling the augmentation pipeline to culturally diverse settings, addressing remaining context-preservation challenges, and assessing how these methods generalize across platforms and content types.
@@ Acknowledgment
This work is supported by the Singapore Ministry of Education AcRF Tier 3 Grant (MOE-MOET32022-0001) and the A*STAR Online Trust and Safety (OTS) Research Programme Grant (S24T2TS011). We are deeply grateful to Rongxin Ouyang for his contributions to the label augmentation component of this work (Ouyang et al. 2025) and for the many insightful discussions on bias and equality. We also thank Insyirah Mujtahid, Shaz Furniturewala, and Jingwei Gao for their valuable input and assistance at various stages of this work.